# Performance in solving the Hermitian and pseudo-Hermitian Bethe-Salpeter equation with the Yambo code

### v2

@@ 4 Conclusions
In conclusion, solvers based on external libraries can be reliably used for the Bethe-Salpeter equation (BSE). The situation is somewhat different from other cases, such as the DFT eigenvalue problem, where the solver is one of the most demanding steps. Accordingly, for the DFT case, it might be better to have DFT-specific and more efficient iterative solvers directly inside the code. For the BSE case, the building of the matrix remains, in many cases, the most demanding step on which code developers can focus. Therefore, opting for library-based solvers is a valuable solution, because it allows us to overcome the “solver barrier”, which appears for very large matrices, especially when extraction of the eigenvectors is needed. Beyond the libraries considered in this work, we are also interfacing the Yambo code with the Magma library [49], and with the cuSolver and cuSolverMP/cuSolverMG libraries [50]. The biggest advantage for BSE code developers in using libraries is that they outsource the implementation, MPI distribution, and GPU (or other architecture) porting. This approach is well justified by the performance results presented in this work.
@@ Acknowledgements
Innovation Study ISOLV-BSE has received funding through the Inno4scale project, which is funded by the European High-Performance Computing Joint Undertaking (JU) under Grant Agreement No 101118139. The JU receives support from the European Union’s Horizon Europe Programme. DS acknowledges funding from the MaX “MAterials design at the eXascale” project, co-funded by the European High Performance Computing joint Undertaking (JU) and participating countries (Grant Agreement No. 101093374). DS also acknowledges funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie grant agreement 101118915 (project TIMES). DS and PM acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support, through the project IsCc6_ISOL-RUN. BM, FA, ER, and JER. were partly supported by grant PID2022-139568NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. BM was also supported by Universitat Politècnica de València in its PAID-01-23 programme. NM and LW acknowledge the use of HPC facilities of the University of Luxembourg [51].

### v3

@@ 4 Conclusions
In conclusion, solvers based on external libraries can be reliably used for the Bethe-Salpeter equation (BSE). The situation is somewhat different from other cases, such as the DFT eigenvalue problem, where the solver is one of the most demanding steps. Accordingly, for the DFT case, it might be better to have DFT-specific and more efficient iterative solvers directly inside the code. For the BSE case, the building of the matrix remains, in many cases, the most demanding step on which code developers can focus. Therefore, opting for library-based solvers is a valuable solution, because it allows us to overcome the “solver barrier”, which appears for very large matrices, especially when extraction of the eigenvectors is needed. Beyond the libraries considered in this work, we are also interfacing the Yambo code with the Magma library [magma_paper], and with the cuSolver and cuSolverMP/cuSolverMG libraries [cusolver]. The biggest advantage for BSE code developers in using libraries is that they outsource the implementation, MPI distribution, and GPU (or other architecture) porting. This approach is well justified by the performance results presented in this work.
@@ Acknowledgements
Innovation Study ISOLV-BSE has received funding through the Inno4scale project, which is funded by the European High-Performance Computing Joint Undertaking (JU) under Grant Agreement No 101118139. The JU receives support from the European Union’s Horizon Europe Programme. DS acknowledges funding from the MaX “MAterials design at the eXascale” project, co-funded by the European High Performance Computing joint Undertaking (JU) and participating countries (Grant Agreement No. 101093374). DS also acknowledges funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie grant agreement 101118915 (project TIMES). DS and PM acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support, through the project IsCc6_ISOL-RUN. BM, FA, ER, and JER. were partly supported by grant PID2022-139568NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. BM was also supported by Universitat Politècnica de València in its PAID-01-23 programme. NM and LW acknowledge the use of HPC facilities of the University of Luxembourg [VCPKVO_HPCCT22].